OpenStudy (anonymous):

Can this be simplified further? (c^2 - 8c) / c

OpenStudy (anonymous):

yes, you can either factor the numerator as \[\frac{c^2-8c}{c}=\frac{c(c-8)}{c}\] and then cancel the \(c\) or else divide each term in the numerator by \(c\) in either case you will get \(c-8\)

OpenStudy (anonymous):

factor out c (c^2 - 8c)/c c(c-8)/c c-8

OpenStudy (anonymous):

Thanks!